The perimeter of a rectangle is 20x-14. The length is 7x+1. Write a polynomial in x in the simplest form for the width. The answer W =?

Answer by tommyt3rd(5050)

P=2(l+w)
P=20x-14
l=7x+1

20x-14=2(7x+1+w)
D 2:

10x-7=7x+1+w

w=3x-8


:)